# Filling of incomplete sinograms from sparse PET detector configurations using a residual U-Net

### v1

@@ 1 Introduction
Positron Emission Tomography (PET) imaging has become a cornerstone in medical diagnostics. Technology-driven developments such as iterative image reconstruction and increased computational power have improved performance and paved the way for constructing long axial field-of-view (FOV) ”total” body PET systems. Current commercial models like the United Imaging uExplorerTM total-body PET scanner and the Siemens Biograph Vision QuadraTM nearly total-body PET scanner, offering axial FOVs of up to 194 cm and 106 cm, respectively, are examples of this latest technological leap [1, 2]. The updated PennPET Explorer further exemplifies this trend towards extended FOV [3, 4, 5]. These advancements promise benefits like lower radioactive doses, shorter scan times, and reduced image noise. However, they come at a significant cost, primarily due to the increased number of detectors required [6, 7, 8, 9]. Therefore, the traditional cylindrical configuration of PET scanners, characterised by densely packed photodetectors, presents substantial financial limitations in the transition to extended FOV PET imaging.
@@ 2.1 Data description and processing
The PET dataset originates from a clinical trial, MORRIS (NCT02379039), approved by the Regional Ethical Review Board, Umeå (dnr 2015/117-31). To evaluate the proposed method, 8 different pelvis scans acquired without time-of-flight on a GE Signa PET/magnetic resonance (MR) imaging system were used. The original dataset, therefore, contained information from all detector positions. To simulate a sparse PET configuration, a subsampling scheme cancelling half of the detectors in a chessboard pattern was adopted (Figure 1). The chessboard pattern was created by cancelling single crystals units, i.e. $1\times 1$ crystals in each chessboard square, resulting in gaps of size $4\times 5.3$mm. Detector removal was performed by discarding coincidences from the original list mode data file and binning the new list mode file into 1981 2D sinograms per scan (the scanner’s standard sinogram format). The resulting count loss corresponds to deleting approximately 75% of the lines of response. In the sparse configuration, we evaluated cancelling out crystal units corresponding to both black and complimentary white positions as shown in Figure 1. The modified dataset contained $8\times 2\times 1981=31696$ incomplete 2D sinogram slices, counting 8 individual scans, chessboard and complimentary chessboard patterns, and all 1981 direct and oblique planes.
@@
Figure 2: Visualisation of the sinogram distortion due to the sparse PET chessboard configuration. (a) Pixel-wise correlation between the original and distorted sinograms; figure shows random sample of $10^{5}$ pixels from the different scans (different colours for different scans). The distorted pixels follow two separate distributions (marked by arrows) – they are either zero (i) or of lower intensity than the original (ii). The ideal fit is illustrated by a solid line, and the overall total fit to the original pixels is illustrated by a dashed line. (b) shows the sinogram distortion pattern with zeroed out pixels in direct planes and cross planes with ring difference $>1$. (c) shows the sinogram distortion pattern with low-intensity pixels in summed cross planes with ring difference 1.
@@ 2.4 Quantification
All data processing steps, including list mode alterations, binning into sinograms, and image reconstructions, were performed with GE Healthcare’s PET research toolbox Duetto (v02.18). Following the clinical protocol, we performed image reconstruction using the 3D ordered subset expectation maximisation algorithm with 28 subsets, 2 iterations, and a 5mm Gaussian postfilter.
For a robust and quantitative comparison, two statistical tests were chosen based on the nature of the data and the aspects of the prediction methods being evaluated. Fisher’s Z-transformation was used to compare pixel-wise correlations between original and predicted sinograms. The Mann-Whitney U test was employed to evaluate significant difference between evaluated measures in both sinograms and reconstructed images.
@@ 3 Results
This section will refer to the ground truth original scans as ”original” and the sparse PET configuration-modified data as ”distorted”. Predictions resulting from the sinogram restoration network are referred to as the ”restored” sinograms, as opposed to the ”interpolated” sinograms that served as a baseline method to fill in the missing data.
@@ 3.1 Sinogram restoration
The interpolated sinograms also present low MAE (Figure 6(b)) and a strong positive correlation with the original counts (Figure 5(a)). However, the correlation is significantly lower compared to the restored predictions (${p<0.001}$). The interpolation method fails to recover the lower counts in the diagonal cross planes of the distorted data, leading to distinctive underrepresentation in the added cross-plane sinograms for ring differences equal to one (Figure 4(b)). The shortcoming of the interpolation method is evident, with inferior values of the SSIM and MAE (${p<0.001}$), as seen in Figure 6.
@@ 3.2 Reconstructed images
The effects of sinogram restoration are also visible as improved image quality of reconstructed images. Figure 7 show images reconstructed from original, distorted, interpolated, and restored sinograms. The smooth nature of the restored sinograms leads to reconstructed images with smoother background texture and a lack of some finer details.
Figure 7: Visual comparison of reconstructed image slices based on original sinograms (Original), sinograms from the sparse chessboard configuration (Distorted), and predicted sinograms from interpolation (Interpolated) and from the restoration network (Restored). (a) and (b) show two different image slices from two different scans in the test set, including corresponding difference images and histograms of pixel differences below each image slice.
The underestimation in interpolated sinogram counts seen in Figure 4(b) has a pronounced impact on reconstruction, resulting in image slices with lower intensity (Figures 7(a) and 8(a)). Moreover, the interpolated sinograms contribute to occasional spurious features in the reconstructed images, appearing as overestimated intensities, visible in Figures 7(b) and 8(a). This tendency is reflected in the histograms in Figure 7, where the interpolated images exhibit a broader and more uneven distribution of pixel differences relative to the original images. In contrast, the restored images show a narrower error distribution, indicating closer overall agreement with the original, despite the inherent smoothing. The visual comparison is also appreciated in the quantitative measures in Figure 9, with superior values of SSIM and MAE for the restored images (${p<0.001}$).
@@ 4 Discussion
The restoration network successfully recovered distorted sinograms. However, the smooth nature of the restored sinograms and the consequential absence of some finer details in reconstructed images help suppress spurious noise but but at the cost of lower resolution, limiting clinical usability at this point. Although small high-activity regions are detected in the reconstructed images, the model demonstrates a lack of precision in correctly quantifying these areas.
@@
The proposed sparse chessboard configuration is designed to offer low-cost or cost-effective clinical PET scanners with the same axial FOV as a conventional system but with only 50% of the original detector materials, thereby halving the cost of the detectors. We further speculate that the proposed sparse geometry could aleternatively be used to allow the development of a two-fold extension of the axial FOV of conventional compact ring configurations with the same number of detectors. An extended axial FOV would, for instance, allow simultaneous imaging of distant organs such as the brain and heart, as in the total body PET systems available on market. This enhanced coverage enables both static and dynamic whole-body imaging, supporting parametric PET – a challenge with standard systems due to axial limitations. A sparse long axial FOV PET scanner with 50% reduction in detectors, resulting in only 25% of the initial LORs retained, would still generate high sensitivity, but 25% of that of the ultra-high sensitivity associated with total body scanners with complete detector coverage. However, our solution would offer the possibility for dynamic whole-body scans at a reasonable cost and sensitivity. Combining the two approaches of extended axial FOV and reduced cost compared to compact designs results in a medium-cost extended FOV, which offers most of the advantages of a full total body PET but positioning itself between the current conventional PET and extended FOV systems in terms of investments. Our method’s usefulness in reconstructing sparse detector arrays from long axial FOV scanners would, of course, need proper validation in future studies.
@@
Future work should include an investigation of incorporating sharpness to the U-Net structure to mitigate the smoothing effect inherent in neural networks. It would also be interesting to investigate this specific model’s capacity regarding possible data losses for clinically approved image quality and if there is an optimal sampling pattern. One might consider an adaptive sparsity pattern, e.g., tighter sampling at the scanner centre and more sparse sampling at the scanner ends. Another necessary step is to explore the integration of sinogram re-normalisation to account for the missing lines of response, particularly in the cross-planes. In this study, we did not explicitly apply re-normalisation. However, our results show that the network is capable of restoring both the missing signals in the direct planes and the attenuated signals in the cross-planes without a separate re-normalisation. This suggests that the model may already be learning an implicit form of normalisation. Nonetheless, applying explicit re-normalisation as a preprocessing step may reduce the network’s burden and improve restoration accuracy – a possibility that warrants further investigation.
@@
Furthermore, the current model was trained on one specific detector pattern at a time. To increase model generalisability, the training process could be extended to include several pattern sizes. Our model could also be combined with parallax correction. In the case of missing LORs due to the missing detector elements, such a correction scheme would also give the added benefit of shifting some of the measured counts to empty detector positions, thereby assisting the estimation of missing counts. Another alternative is to extend our model to include time-of-flight (TOF) information. We expect that including the TOF timing resolution would improve lesion detectability compared to the non-TOF setting and increase the number of 2D sinograms available for training. An additional future direction would be to evaluate whether combining our method with, for instance, bed motion could enhance the sinogram restoration, which allows measuring the same patient position with different detector configurations.
@@ Acknowledgments
This work is supported by the Swedish Research Council (340-2013-5342).

### v2

@@ 1 Introduction
Positron Emission Tomography (PET) imaging has become a cornerstone in medical diagnostics. Technology-driven developments such as iterative image reconstruction and increased computational power have improved performance and paved the way for constructing long axial field-of-view (FOV) "total" body PET systems. Current commercial models like the United Imaging uExplorerTM total-body PET scanner and the Siemens Biograph Vision QuadraTM nearly total-body PET scanner, offering axial FOVs of up to 194 cm and 106 cm, respectively, are examples of this latest technological leap [1, 2]. The updated PennPET Explorer further exemplifies this trend towards extended FOV [3, 4, 5]. These advancements promise benefits like lower radioactive doses, shorter scan times, and reduced image noise. However, they come at a significant cost, primarily due to the increased number of detectors required [6, 7, 8, 9]. Therefore, the traditional cylindrical configuration of PET scanners, characterised by densely packed photodetectors, presents substantial financial limitations in the transition to extended FOV PET imaging.
@@ 2.1 Data description and processing
The PET dataset originates from a clinical trial, MORRIS (NCT02379039), approved by the Regional Ethical Review Board, Umeå (dnr 2015/117-31). To evaluate the proposed method, 8 different pelvis scans acquired without time-of-flight on a GE Signa PET/magnetic resonance (MR) imaging system were used. The original dataset, therefore, contained information from all detector positions. To simulate a sparse PET configuration, a subsampling scheme cancelling half of the detectors in a chessboard pattern was adopted (Figure 1). The chessboard pattern was created by cancelling single crystals units, i.e. $1\times 1$ crystals in each chessboard square, resulting in gaps of size $4\times 5.3$mm. Detector removal was performed by discarding coincidences from the original list mode data file and binning the new list mode file into 1981 2D sinograms per scan (the scanner's standard sinogram format). The resulting count loss corresponds to deleting approximately 75% of the lines of response. In the sparse configuration, we evaluated cancelling out crystal units corresponding to both black and complimentary white positions as shown in Figure 1. The modified dataset contained $8\times 2\times 1981=31696$ incomplete 2D sinogram slices, counting 8 individual scans, chessboard and complimentary chessboard patterns, and all 1981 direct and oblique planes.
@@
Figure 2: Visualisation of the sinogram distortion due to the sparse PET chessboard configuration. (a) Pixel-wise correlation between the original and distorted sinograms; figure shows random sample of $10^{5}$ pixels from the different scans (different colours for different scans). The distorted pixels follow two separate distributions (marked by arrows) – they are either zero (i) or of lower intensity than the original (ii). The ideal fit is illustrated by a solid line, and the overall total fit to the original pixels is illustrated by a dashed line. (b) shows the sinogram distortion pattern with zeroed out pixels in direct planes and cross planes with ring difference >1>1. (c) shows the sinogram distortion pattern with low-intensity pixels in summed cross planes with ring difference 1.
@@ 2.3 Network training
The model was developed using the Keras modules of TensorFlow and trained on a high-performance computer equipped with 48 CPU cores and 2 NVIDIA A100 GPUs. The data was converted into the TFRecords format, and the loading pipeline was optimized, so each epoch took about 36 seconds. The minimum, maximum, and average training time (number of epochs) were 73 minutes, 122 minutes, and 98 minutes (109, 194, 149 epochs), respectively. Once the model was trained, the runtime to infer the restored sinograms was approximately of 20 seconds per patient. The same operation took approximately 2 and a half minutes on an Apple MacBook Pro M1 (2021), with a much more modest hardware configuration which is comparable to the ones found within typical clinical settings.
@@ 2.4 Quantification
To complement these global metrics, we also evaluated background variability (BV), defined as the coefficient of variation of ROI mean values, expressed as a percentage:
$BV=\frac{\sigma_{ROI}}{\mu_{ROI}}\cdot 100$
where $\mu_{ROI}$ is the average ROI mean value and $\sigma_{ROI}$ is the standard deviation across ROI mean values. In this study, we report a relative background variability (rBV), defined as the ratio of BV in the restored images to BV in the corresponding original images.
All data processing steps, including list mode alterations, binning into sinograms, and image reconstructions, were performed with GE Healthcare's PET research toolbox Duetto (v02.18). Following the clinical protocol, we performed image reconstruction to 192×192×89 voxels, 60 cm radial and 25 cm axial FOV via the 3D ordered subset expectation maximisation algorithm with 28 subsets, 2 iterations, and a 5mm Gaussian postfilter. The reconstruction used the default normalisation file and corrections for attenuation, scatter and randoms.
For a robust and quantitative comparison, two statistical tests were chosen based on the nature of the data and the aspects of the prediction methods being evaluated. Fisher's Z-transformation was used to compare pixel-wise correlations between original and predicted sinograms. The Mann-Whitney U test was employed to evaluate significant difference between evaluated measures in both sinograms and reconstructed images.
@@ 3 Results
This section will refer to the ground truth original scans as "original" and the sparse PET configuration-modified data as "distorted". Predictions resulting from the sinogram restoration network are referred to as the "restored" sinograms, as opposed to the "interpolated" sinograms that served as a baseline method to fill in the missing data.
@@ 3.1 Sinogram restoration
The interpolated sinograms also present low MAE (Figure 6(b)) and a strong positive correlation with the original counts (Figure 5(a)). However, the correlation is significantly lower compared to the restored predictions (p<0.001{p<0.001}). The interpolation method fails to recover the lower counts in the diagonal cross planes of the distorted data, leading to distinctive underrepresentation in the added cross-plane sinograms for ring differences equal to one (Figure 4(b)). The shortcoming of the interpolation method is evident, with inferior values of the SSIM and MAE (p<0.001{p<0.001}), as seen in Figure 6.
The interested reader is referred to supplementary Figure S.1 to inspect the same visual comparison of summed cross-plane sinogram slices with ring difference greater than 1.
@@ 3.2 Reconstructed images
[…]
Figure 7: Visual comparison of reconstructed image slices based on original sinograms (Original), sinograms from the sparse chessboard configuration (Distorted), and predicted sinograms from interpolation (Interpolated) and from the restoration network (Restored). (a) and (b) show two different image slices from two different scans in the test set, with reconstructed images on the top vertical panel, corresponding difference images on the middle panel (colour bar in Bq/ml), and corresponding histograms of pixel differences in the bottom panel ($x$-axis in Bq/ml).
The underestimation in interpolated sinogram counts seen in Figure 4(b) has a pronounced impact on reconstruction, resulting in image slices with lower intensity (Figures 7(a) and 8(a)). Moreover, the interpolated sinograms contribute to occasional spurious features in the reconstructed images, appearing as overestimated intensities, visible in Figures 7(b) and 8(a). This tendency is reflected in the histograms in Figure 7, where the interpolated images exhibit a broader and more uneven distribution of pixel differences relative to the original images. In contrast, the restored images show a narrower error distribution, indicating closer overall agreement with the original, despite the inherent smoothing. The visual comparison is also appreciated in the quantitative measures in Figure 9, with superior values of SSIM and MAE for the restored images (p<0.001{p<0.001}).
Image slices from the same scans of Figure 7 but reconstructed in coronal view can be found in supplementary Figure S.2.
@@ 4 Discussion
The restoration network successfully recovered distorted sinograms. However, the smooth nature of the restored sinograms causes excess smoothing also in the reconstructed images, limiting clinical usability at this point. Although small high-contrast regions are detected in the images, the model demonstrates a lack of precision in correctly quantifying these areas. The smoothing effect seems to have a direct impact on the contrast recovery, which was evident in the reconstructed image in Figure 7 where some high-contrast areas are reduced in size and some are distorted in their shape. It is also evident in the pixel-wise contrast comparison in Figure 8(b), where we see an indication of high-contrast pixels as the main drivers to the deviation from a perfect fit to the original pixels.
@@
The proposed sparse chessboard configuration is designed to explore the potential for lower-cost or more cost-effective clinical PET scanners, aiming for the same axial FOV as conventional scanners but with only 50% of the original detector materials. We further speculate that the proposed sparse geometry could alternatively be used to allow the development of a two-fold extension of the axial FOV of conventional compact ring configurations with the same number of detectors. An extended axial FOV would, for instance, allow simultaneous imaging of distant organs such as the brain and heart, as in the total body PET systems available on the market. This enhanced coverage enables both static and dynamic whole-body imaging, supporting parametric PET – a challenge with standard systems due to axial limitations. A sparse long axial FOV PET scanner with 50% fewer detectors, retaining only 25% of the initial LORs, would inevitably have substantially lower sensitivity than full-coverage total-body scanners. However, its sensitivity would remain comparable to that of earlier-generation PET systems, which have more limited axial coverage. From this perspective, such a system could still enable dynamic whole-body imaging at a reasonable cost and acceptable sensitivity. Combining the two approaches of extended axial FOV and reduced detector cost compared to compact designs results in a medium-cost extended FOV, which offers many of the advantages of a full total body PET but positioning itself between the current conventional PET and extended FOV systems in terms of investments. We note, however, that while the 1×1 sparsity pattern performed best in our evaluation, such a configuration may not be the most practical or cost-effective from a manufacturing perspective. The present work should therefore be regarded as a proof-of-concept from a deep learning standpoint, with practical implementation requiring further study.
@@
Future work should include an investigation of incorporating sharpness to the U-Net structure to mitigate the smoothing effect inherent in neural networks. It would also be interesting to investigate this specific model's capacity regarding possible data losses for clinically approved image quality and if there is an optimal sampling pattern. One might consider an adaptive sparsity pattern, e.g., tighter sampling at the scanner centre and more sparse sampling at the scanner ends. Another necessary step is to explore the integration of sinogram re-normalisation to account for the missing lines of response, particularly in the cross-planes. In this study, we did not explicitly apply re-normalisation. However, our results show that the network is capable of restoring both the missing signals in the direct planes and the attenuated signals in the cross-planes without a separate re-normalisation. This suggests that the model can, to some extent, compensate for the effects of missing and attenuated signals without explicit need for re-normalisation of the sinograms. Nonetheless, applying explicit re-normalisation as a preprocessing step may reduce the network's burden and improve restoration accuracy – a possibility that warrants further investigation.
@@
Furthermore, the dataset was limited in the sense that it only included one anatomical region and only left one patient for validation. This was a main motivation for training the model on individual sinogram slices, which comes at a possible trade-off of introduced bias in the results. The current model was furthermore trained on one specific detector pattern at a time. To increase model generalisability, the dataset should be extended to multiple anatomical regions, and the training process could be extended to include several pattern sizes. Our model could also be combined with parallax correction. In the case of missing LORs due to the missing detector elements, such a correction scheme would also give the added benefit of shifting some of the measured counts to empty detector positions, thereby assisting the estimation of missing counts. Another alternative is to extend our model to include time-of-flight (TOF) information. We expect that including the TOF timing resolution would improve lesion detectability compared to the non-TOF setting due to enhanced signal-to-noise, mainly due to the increase the number of 2D sinograms available for training. We do not expect the reported blurring to go away since the same detector elements are still missing. An additional future direction would be to evaluate whether combining our method with, for instance, bed motion could enhance sinogram restoration, since bed motion effectively provides multiple slightly shifted sampling configurations of the same patient position.
@@ Acknowledgments
The authors acknowledge the use of high-performance computing resources provided by High Performance Computing Center North (HPC2N), which supported the computational work for this study. This work is supported by the Swedish Research Council (340-2013-5342).